# A Survey on the Applications of Zero-Knowledge Proofs

### v1

@@ II-B2 Arithmetization: From circuits to matrices
Arithmetization is the process of encoding the wiring and gate operations of a digital circuit or model of computation into a mathematical representation. In the context of zero-knowledge proofs, this often involves translating the circuit into a format that is suitable for the generation of a proof that the circuit’s execution satisfies certain conditions without revealing the specifics of the inputs or internal states. For brevity, we will focus on Rank-1 Constraint Systems (R1CS), which serve as a common arithmetization strategy for many zk-SNARKs. Some alternative arithmetization methods, namely Plonk-ish arithmetization, algebraic intermediate representations (AIR), and circuit constraint systems (CCS) have gained popularity in recent years as well [15, 16].
An R1CS is a way to represent a computation in the form of constraints that are linear equations. These constraints can be checked efficiently and are the foundation of various zero-knowledge proof systems. In an R1CS, each constraint is an equation of the form:
$\mathbf{A}\cdot\mathbf{w}\circ\mathbf{B}\cdot\mathbf{w}=\mathbf{C}\cdot\mathbf{w}$
@@
where $\mathbf{A}$, $\mathbf{B}$, and $\mathbf{C}$ are matrices, $\mathbf{w}$ is a vector of variables that correspond to the transcript of a digital circuit’s execution, and $\circ$ denotes the Hadamard (element-wise) product.
As a mental model, one can think of R1CS as a system that verifies all parts of a program by checking individual, linear parts of its computation, ensuring that all constraints are satisfied for the program to be considered correct.
The operations performed by our example circuit can be encoded into an R1CS by expressing each operation as a set of linear constraints. In R1CS, the constraints take the form of equations where the left and right sides must multiply to give the output. For our circuit, we introduce variables and intermediate ’witness’ variables to represent the computation steps. Here is how each operation in the circuit translates into constraints:
@@
    Multiply $a$ by itself to get $a^{2}$ results in the constraint $a\times a-a^{2}=0$.
@@
    Multiply $b$ by itself to get $b^{2}$ results in the constraint $b\times b-b^{2}=0$.
@@
    Multiply $c$ by itself to get $c^{2}$ results in the constraint $c\times c-c^{2}=0$.
  4. 4.
    Multiply $a^{2}$ by $b^{2}$ to get $a^{2}\times b^{2}$ results in the constraint $a^{2}\times b^{2}-a^{2}b^{2}=0$.
  5. 5.
    Multiply $b^{2}$ by $c^{2}$ to get $b^{2}\times c^{2}$ results in the constraint $b^{2}\times c^{2}-b^{2}c^{2}=0$.
  6. 6.
    Add $a^{2}\times b^{2}$ and $b^{2}\times c^{2}$ to get the final result leads to the constraint $ab^{2}+bc^{2}-result=0$.
To encode these constraints into matrices $\mathbf{A}$, $\mathbf{B}$, and $\mathbf{C}$, we introduce the following variables: $w_{1}=a$, $w_{2}=b$, $w_{3}=c$, $w_{4}=a^{2}$, $w_{5}=b^{2}$, $w_{6}=c^{2}$, $w_{7}=a^{2}b^{2}$, $w_{8}=b^{2}c^{2}$, and $w_{9}=result$. Each matrix will have rows corresponding to constraints and columns corresponding to these variables, which can be further compressed as an array $\mathbf{w}=[1,a,b,c,a^{2},b^{2},c^{2},ab^{2},bc^{2},\text{result}]$.
The vector $\mathbf{w}$, often referred to as the ‘witness’, captures an execution transcript of the circuit. The witness contains the values of all variables and intermediate steps that satisfy all the R1CS constraints for the given inputs. Given that anyone with a witness $\mathbf{w}$ can execute the circuit to verify if $\mathbf{w}$ is valid, a correct execution transcript $\mathbf{w}$ can be viewed as a proof that the computation was carried out correctly and that all of the R1CS’s constraints are satisfied. If, at any point, an attacker obtained access to the R1CS or witness vector, access to the inputs, intermediate values, and outputs of the computation would be leaked.
Up to this point, there is no way to prove, in a publicly-verifiable way, that the computation was performed correctly without allowing others to re-run the computation themselves. The next section, however, converts the R1CS into a form that allows for maintaining privacy while allowing for public-verifiability and a sublinear verifier runtime.
II-B3 Backends: From matrices to polynomials
The backend phase of the SNARK lifecycle involves transforming the Rank-1 Constraint System (R1CS) matrices into a set of polynomial equations that form a Quadratic Arithmetic Program (QAP). This transformation is key to creating a SNARK, as it allows for the efficient verification of the computations represented by the R1CS. Succinctness, a property of SNARKs integral to a sub-linear verifier runtime, is achieved by encoding the matrix operations from the R1CS into polynomial equations. This more succinctly represents the original circuit’s execution trace without sacrificing the soundness guarantees of the underlying argument.
Mathematically, a QAP is defined as a set of polynomial equations derived from the R1CS [17]. For matrices $\mathbf{A},\mathbf{B},\mathbf{C}$ of size $m\times n$, where $m$ is the number of constraints and $n$ is the number of variables, we define the QAP as follows: Let $\{A_{i}(x)\}_{i=1}^{m}$, $\{B_{i}(x)\}_{i=1}^{m}$, and $\{C_{i}(x)\}_{i=1}^{m}$ be sets of polynomials where each $A_{i}(x),B_{i}(x),C_{i}(x)$ is a polynomial corresponding to the $i^{th}$ row of $\mathbf{A},\mathbf{B},\mathbf{C}$ respectively. Then, the QAP consists of the set of equations $\{A_{i}(x)\cdot B_{i}(x)-C_{i}(x)=h(x)\cdot Z(x)\}_{i=1}^{m}$, where $h(x)$ is the polynomial that represents a solution to the R1CS, and $Z(x)$ is a verifier-chosen, fixed polynomial that evaluates to zero at $\{x_{1},x_{2},\ldots,x_{m}\}$, the set of distinct points associated with each constraint.
The QAP allows for the succinct representation of the R1CS and enables the prover to compute a single proof, which is a small set of polynomial evaluations. This proof attests that there exists a polynomial $h(x)$ that satisfies the QAP equation, implying that the prover knows a witness $\mathbf{w}$ that satisfies the R1CS.
The verifier, given the proof and the public polynomials $A(x)$, $B(x)$, and $C(x)$, can efficiently check the QAP verification equation without having to perform the full set of R1CS computations. This results in a verification process with a significant speedup over computing the function represented by the circuit directly.
In summary, the conversion from matrices to polynomials and the creation of a QAP forms the bridge between the arithmetic circuit and the final SNARK. This process ensures that a valid execution of the circuit, one that satisfies all the constraints, is encoded into a format that supports the creation of a compact, easily verifiable proof. The integrity of this transformation is fundamental to the trustworthiness and security of the SNARK.
II-C Conjoining Information Theory and Cryptography
Interactive proofs (IPs), when not paired with any cryptographic protocol, provide unconditional soundness, where the validity of a statement can only be confirmed if it is indeed true, independent of computational assumptions. Integrating an IP protocol with a cryptographic gadget, however, enhances both the security and other properties of the resulting SNARK. This cryptographic gadget, often referred to as a polynomial commitment scheme (PCS) [18], increases security assumptions but is also a large part of what makes SNARKs practical today through the succinctness property and an improved verifier runtime. While PCSs have become a necessary component in the large taxonomy of SNARK designs today, the necessity for private randomness in PCSs for cryptographic security complicates their integration with an IP protocol. Trusted ceremonies, which are critical for generating this randomness securely and ensuring that no single participant can compromise the setup, thus need to be executed before the SNARK enters production. If a trusted ceremony is executed incorrectly, false proofs will be able to be created and the protocol will be insecure – this happened in 2016 with Zcash’s Sapling ceremony and another ceremony was later ran to replace it [19]. Lastly, to ensure that SNARKs are publicly-verifiable, so that collusion between the prover and verifier is not possible, the SNARK must be made non-interactive. SNARKs achieve non-interactivity through the Fiat-Shamir heuristic [20], which substitutes the verifier’s random challenges with outputs from a collision-resistant hash function. This modification not only simplifies the verification process but also makes the proofs publicly verifiable, enhancing both the practicality and security of SNARKs in broad applications.
III ZKP Software Tools and Platforms
In this section, we explore the underlying infrastructure and tooling currently available to produce ZKP by first examining zero-knowledge virtual machines (zkVM), domain specific languages (DSLs), and supporting libraries, frameworks, and protocols. These core pieces of the ZKP space allow for statements we are familiar with to be translated into an arbitrary computation that is then represented as a verifiable ZK circuit in an efficient manner.
At a broad scale, the process can be split into two sections: the frontend and the backend. Given a program, the process begins with the frontend. The frontend’s purpose is to take in a program and transform it into a format that a proving system discussed above can interpret and produce a proof from. The backend handles the proof generation process and does most of the heavy lifting in the process.
After that, the intermediate representation is translated into a constraint system and passed to the backend. The backend begins by putting the circuit representation of the program from the artihmetization scheme into a proving system such as Groth16 [21] or Bulletproofs [22]. Finally, a proof of execution is produced and the user can succinctly verify its correctness.
III-A Zero-Knowledge Virtual Machines
III-A1 Motivation and Definition
With the increased dependency on distributed computing in industries such as blockchain and cloud computing, the importance of private computing at a large scale grew exponentially. ZKPs offer a solution for providing an environment that performs computations in a trustworthy manner through a zero-knowledge virtual machine (zkVM). A zkVM is a virtual machine (VM) designed to generate an efficiently verifiable proof for executing an arbitrary computation while preserving the privacy of the program and its data in zero knowledge. The zkVM acts as a programmable ZK circuit that implements a VM, receives public and private inputs, and creates a certificate of valid execution. Similar to a generalized virtual machine, a zkVM maintains its own virtualized components, such as memory management, instruction scheduling, error handling, and others, using a privacy-focused approach with the help of a ZKP.
The motivation for building zkVMs stems from the desire to reduce the complexity of creating a ZKP. Previously, to produce a ZKP, a developer needed to have a strong background in cryptography, develop a circuit representation of their program, set up a backend for each proof, and have a large amount of computing power. To abstract away these steps from the developer, zero-knowledge virtual machines were developed with these features in mind.
III-A2 Methodology
zkVMMain MemoryOperating SystemRegistersCPUBackendProofProgram OutputProgram CodeWitnessCommitments
Figure 3: General zkVM Architecture
The first step of most zkVMs is to receive an input program from a higher-level language such as C++, Circom [23], Rust, or others. The program is then compiled to bytecode that the VM can translate to an instruction set architecture (ISA) that is generalized, like RISCV, or specialized, such as Miden Assembly [24]. These ISAs are minimal and optimized for cryptographic operations such as hashing to optimize the performance of the zkVM.
Depending on the proving system, the setup phase begins after the program has been interpreted and translated to bytecode. This step involves tasks such as concealing private inputs, initializing a polynomial commitment scheme, key generation, and others. A set of constraints is placed on the program and the witness to enforce computational integrity in the computation trace. Building on what was mentioned in Section 2, constraints ensure the zkVM execution trace follows the guidelines outlined in the ISA. Every constraint is a polynomial for each statement inside the execution trace.
To check the validity of each instruction in the execution trace, a verifier can open any instruction in the zkVM to verify it was computed correctly. The prover defines a polynomial for each statement that the verifier can use the polynomial commitment scheme to verify. The verifier receives this polynomial and responds with a random value for evaluating the prover’s polynomial. If the verifier concludes that the output is correct, then the statement in question is also considered correct.
The zkVM distinguishes itself from writing a specialized circuit for computations due to its high optimization level and reduction of complexity from designing a circuit. In a ZK computation, three common overheads must be taken into account:
  1. 1.
    Range Checks: A range check is a way to prove that an element in a finite field falls within a specific interval [a, b]. Due to the variance in fields, the importance of correct bit-sized integers in these fields, and the intensive operations throughout them, it is important to verify the inputs are valid. This is done using a range proof to verify that a given integer fits the necessary criteria, such as fitting into 32 bits without overflowing.
  2. 2.
    Bitwise operations: Nearly every computation in a ZK context needs some bitwise operation for tasks such as hashing. On standard computer hardware, these operations are highly optimized for numbers of standard sizes rather than for large elements in a finite field. In a ZK context, every bit inside of a field element must be proved, the decomposition of a field element into its bits must be proved, and finally, the result of a bitwise operation must be aggregated from bits into another field operation. With current hardware that is not optimized for ZK computation, these can incur large overheads in common operations.
  3. 3.
    Hashing. Traditional primitives use existing complex optimizations for bitwise operations as mentioned in 2); however, these bitwise operations take more time because of the above reasoning. Fundamental hash functions such as SHA256 become obsolete on zkVMs unless improvements are made. Considering this, many zkVMs use rivaling primitives such as Rescue or Poseidon or use an entirely different processor to handle the hashing.
By prioritizing efficiency in cryptographic protocols that leverage zero-knowledge proofs, zkVM enhances the performance of applications involved in secure data exchange, anonymous transactions, and confidential smart contracts. This optimization ensures that cryptographic operations, integral to privacy-preserving technologies, are executed with heightened speed and effectiveness.
20132015201720192021202220232024TinyRAMHawkZincRISC ZeroCairoVMAleoMidenTritonOlaVMPowdrJoltValidaSP1Nexus
Figure 4: Timeline of zkVM Popular Projects
The projects in the timeline are an overview of the prominent zkVMs over the last decade such as TinyRAM [25], Hawk [26], Zinc [27] , RISC Zero [28], CarioVM [29], Leo [30], Miden [24], Triton [31], OlaVM [32] , Powdr [33], and Jolt [34]. Some other zkVMs that are on the rise as of writing this paper are SP1 [35], Nexus [36], and Valida [37].
A project in the space is RISC Zero’s [28] zkVM which allows users to demonstrate the accurate execution of arbitrary Rust and C++ code. This enables the construction of zero-knowledge applications that leverage the variety of existing Rust packages. Functioning as a verifiable computer, the zkVM mimics the operations of a genuine embedded RISC-V micro-processor, thus allowing any program that can compile down to the RISC-V instruction set to run on the zkVM. This emulation not only facilitates the development process for developers but also simplifies the creation of powerful zero-knowledge applications. RISC Zero also mains a ZK Validity Proof System that empowers users to substantiate the validity of Ethereum blocks. The zkVM is equipped to function on different computer systems, program sizes, hardware targets, and runtime/memory requirements.
Starkware [29], the company behind zk-STARKs (Zero Knowledge Scalable Transparent Argument of Knowledge) applications, created CairoVM. CairoVM is an efficient and practical architecture specifically compatible with the STARK proof system. It instantiates a STARK based von-Neumann architecture known for its instruction set that allows support for traditional language features such as conditional branching, function calls, recursion, and its ability to have a deterministic or a nondeterministic nature in aspects like its restrictive memory model. The bytecode of the Cairo zkVM, which are generated by the Cairo assembler, represents the input program from a given user. From there, the Cairo Runner is used to obtain the execution trace of the program. The Cairo Runner is a computer program designed for the execution of a compiled Cairo program that operates in a distinctive manner compared to running a typical computer program. The key disparity arises from Cairo’s support for nondeterministic code. Finally, a STARK prover for the Cairo AIR mentioned in Section 2 is employed to generate a proof for the assertion from the program.
Polygon’s Miden VM [24] is another zkVM built to contribute to a zero-knowledge rollup, which is a modular second layer blockchain built on top of another blockchain such as Ethereum. Miden extends the capabilities of Ethereum by introducing features such as parallel transaction execution and client-side proving. This empowers developers to craft high-throughput and privacy-preserving decentralized applications (dApps) for sectors like DeFi, RWA, and the Metaverse. Popular programming languages like Rust are supported by Miden, which gives developers flexibility and familiarity. Another feature of Miden VM lies in its automatic generation of a STARK-based proof of execution for any program that runs on it. Miden VM operates by consuming programs in the form of a Merkelized Abstract Syntax Tree (MAST), where each node represents a code block in a binary tree structure. Execution begins at the tree’s root, recursively processing each required block based on its semantics. If the execution of any code block encounters failure, the VM halts, preventing the execution of further blocks. The program structure involves a binary tree arrangement, with leaves containing linear instruction sequences and internal nodes determining control flow. As a stack machine, Miden VM utilizes a push-down stack of practically unlimited depth, aligning with the architecture of a STARK. The Miden standard library enhances functionality by providing optimized implementations of commonly-used primitives. This library facilitates the reduction of shared code between parties involved in proving and verifying program execution, achieved by serializing calls to standard library procedures as a fixed 32 bytes, irrespective of the procedure’s size. These design goals collectively contribute to Miden VM’s efficiency, flexibility, and security.
In conclusion, zkVMs offer a broader spectrum of applicability through its support for general and specialized computation. A zkVM is designed to accommodate a wide range of computations unlike some zero-knowledge proof systems. This flexibility permits developers to implement complex algorithms and diverse computations while benefiting from the privacy and security guarantees inherent in zero-knowledge proofs.
III-B Domain Specific Languages
III-B1 Motivation and Definition
A zkDSL (zero-knowledge domain specific language) is a programming language written specifically to provide the programmer a connection between a high-level representation of a program and the low-level intricacies of a ZKP such as writing circuits.

### v2

@@ 2.2.2 Arithmetization: From circuits to matrices
Arithmetization is the process of encoding the wiring and gate operations of a circuit or model of computation into an algebraic representation suitable for proof generation. For brevity, we focus on Rank-1 Constraint Systems (R1CS), a common arithmetization strategy for many zkSNARK constructions. Alternative arithmetizations, including Plonkish arithmetization, algebraic intermediate representations (AIR), and circuit constraint systems (CCS) (see Table 2), have gained popularity in recent years [12, 89].
An R1CS represents a computation as constraints of the form
$(\mathbf{A}\cdot\mathbf{w})\circ(\mathbf{B}\cdot\mathbf{w})=(\mathbf{C}\cdot\mathbf{w}),$
@@
where $\mathbf{A}$, $\mathbf{B}$, and $\mathbf{C}$ are matrices, $\mathbf{w}$ is a vector of variables corresponding to the execution transcript (the witness), and $\circ$ denotes the Hadamard (element-wise) product.
Each gate in the circuit corresponds to one constraint row. For our example, six constraints encode the three squarings ($a\cdot a=a^{2}$, $b\cdot b=b^{2}$, $c\cdot c=c^{2}$), the two cross-multiplications ($a^{2}\cdot b^{2}$, $b^{2}\cdot c^{2}$), and the final addition.
To encode these constraints into matrices $\mathbf{A}$, $\mathbf{B}$, and $\mathbf{C}$, we introduce variables: $w_{1}=a$, $w_{2}=b$, $w_{3}=c$, $w_{4}=a^{2}$, $w_{5}=b^{2}$, $w_{6}=c^{2}$, $w_{7}=a^{2}b^{2}$, $w_{8}=b^{2}c^{2}$, and $w_{9}=\text{result}$. A compact witness vector representation is then
$\mathbf{w}=[1,a,b,c,a^{2},b^{2},c^{2},a^{2}b^{2},b^{2}c^{2},\text{result}].$
The witness $\mathbf{w}$ captures a valid execution transcript of the circuit by assigning values to all intermediate variables so that every constraint is satisfied. If the witness were revealed, it would leak inputs and intermediate values. Up to this point, however, there is still no way to prove correctness succinctly and publicly without allowing others to re-run the computation. The next step converts the R1CS into a polynomial form that enables succinct public verification while preserving privacy in zkSNARK settings.
2.2.3 Backends: From matrices to polynomials
The backend phase transforms the R1CS constraints into polynomial equations, commonly instantiated as a Quadratic Arithmetic Program (QAP) [34]. This transformation enables succinct verification by reducing many constraint checks into a small number of algebraic relations.
For matrices $\mathbf{A},\mathbf{B},\mathbf{C}$ of size $m\times n$, where $m$ is the number of constraints and $n$ is the number of variables, we define polynomials $\{A_{i}(x)\}_{i=1}^{m}$, $\{B_{i}(x)\}_{i=1}^{m}$, and $\{C_{i}(x)\}_{i=1}^{m}$ corresponding to the rows of $\mathbf{A},\mathbf{B},\mathbf{C}$, respectively. The QAP condition can be expressed as:
$A(x)\cdot B(x)-C(x)=h(x)\cdot Z(x),$
where $h(x)$ encodes the satisfaction of the constraints, and $Z(x)$ is a fixed target polynomial that vanishes on the constraint index set.
A prover can produce a compact proof that such an $h(x)$ exists, which implies that the prover knows a witness $\mathbf{w}$ satisfying the original R1CS. A verifier can then check the required polynomial relations efficiently, without re-executing the full computation, yielding a substantial speedup in verification time.
2.3 Conjoining Information Theory and Cryptography
Interactive proofs (IPs), when considered without cryptographic assumptions, can provide information-theoretic soundness. Zero knowledge means that whatever the verifier learns from the interaction could also have been produced by the verifier on its own, without access to the witness, so the protocol leaks nothing beyond the fact that the statement is true. Modern SNARK designs integrate algebraic proof techniques with cryptographic components to achieve succinctness and public verifiability. A common modern design pattern is to combine a polynomial interactive-oracle proof (PIOP/IOP) that reduces computation to a few algebraic checks with a polynomial commitment scheme (PCS), and then apply Fiat–Shamir to remove interaction in the random-oracle model. A common cryptographic component is a polynomial commitment scheme (PCS) [47], which allows the prover to commit to polynomials and later open evaluations in a way that is binding and (often) hiding. This enables efficient verification while preserving soundness under computational assumptions.
Setup assumptions depend on the proof family. In many pairing-based zkSNARK systems, public parameters are generated via a CRS that may depend on secret trapdoor randomness, motivating multi-party ceremonies to reduce trust in any single participant. In transparent families, public parameters can be derived from public randomness, avoiding secret trapdoors.
Finally, non-interactivity is typically achieved either by designs that are non-interactive by construction or by applying the Fiat–Shamir heuristic [30] to suitable interactive protocols, replacing verifier challenges with hash-derived values under a random-oracle-style assumption. This step enables public verifiability and makes the resulting proofs practical for deployment in broad applications.
3 ZKP Software Tools and Platforms
We organize this section using a simple toolchain taxonomy that will be reused in later application comparisons: (i) zkVMs provide a general-purpose execution layer that produces proofs of program execution, (ii) zkDSLs and circuit languages compile high-level intent into circuits or constraints, (iii) libraries and frameworks provide reusable gadgets, curve and field arithmetic, and proof-system implementations, (iv) proving backends implement the concrete proof-system algorithms and polynomial commitment schemes, and (v) hardware accelerators (GPU/FPGA/ASIC) target the dominant kernels such as MSM, FFT/NTT, and hashing.
3.1 Zero-Knowledge Virtual Machines
3.1.1 Motivation and Definition
zkVM is a virtual machine (VM) designed to generate an efficiently verifiable zero-knowledge proof for executing an arbitrary computation while preserving the privacy of the program and its data in zero knowledge. The zkVM acts as a programmable ZK circuit that implements a VM, receives public and private inputs, and creates a certificate of valid execution. Similar to a generalized virtual machine, a zkVM maintains its own virtualized components, such as memory management, instruction scheduling, error handling, and others, using a privacy-focused approach with the help of a ZKP (see Figure 3). zkVMs simplify ZKP creation. Traditionally, developers needed cryptography expertise, circuit representation skills, backend setup for each proof, and significant computing power. zkVMs abstract these complexities, streamlining the process for developers.
3.1.2 Methodology
In Figure 3, the left-side inputs: program code, commitments, and witness represent what the prover supplies to the zkVM, where commitments bind the prover to public artifacts, for example, committed program, committed state, or committed data roots, and the witness contains private execution details. The internal blocks represent the virtualized execution environment (memory, OS/runtime, registers, CPU/ISA interpreter), while the backend represents the proving backend that turns the execution trace into a proof. The right-side outputs are the proof and the program output that a verifier checks. A key takeaway is that zkVM performance is usually dominated by trace generation and the proving backend kernels, not by the conceptual VM blocks themselves. In practice, the cost concentrates in constraint checks for the ISA semantics, memory consistency checks, hashing and commitment openings for trace integrity, and algebraic kernels such as MSM and FFT/NTT depending on the proof family.
zkVMMain MemoryOperating SystemRegistersCPUBackendProgram CodeCommitmentsWitnessProofProgram Output
Figure 3: General zkVM Architecture
The first step of most zkVMs is to receive an input program from a higher-level language such as C++, Circom [20], Rust, or others. The program is then compiled to bytecode that the VM can translate to an instruction set architecture (ISA) that is generalized, like RISCV, or specialized, such as Miden Assembly [80]. These ISAs are minimal and optimized for cryptographic operations such as hashing to optimize the performance of the zkVM. After the program is translated to bytecode, the setup phase involves tasks like concealing private inputs, initializing a polynomial commitment scheme, and key generation. Constraints are then applied to the program and witness to ensure computational integrity, aligning with the ISA. Each constraint corresponds to a polynomial for each statement in the execution trace. To check the validity of each instruction in the execution trace, a verifier can open any instruction in the zkVM to verify it was computed correctly.
The zkVM distinguishes itself from writing a specialized circuit for computations by its high optimization level and reduction of complexity from designing a circuit. In a ZK computation, three common overheads must be taken into account:
@@ 3.1.2 Methodology
    Range Checks: A range check is a way to prove that an element in a finite field falls within a specific interval [a, b].
@@
    Bitwise operations: Nearly every computation in a ZK context needs some bitwise operation for tasks such as hashing, and these operations need to be optimized for elements in finite fields.
@@
    Hashing: Traditional hashes such as SHA256 often incur high overhead in many arithmetizations because they are built from bitwise operations (XOR/AND), rotations, and additions over words, which must be emulated using boolean constraints or non-native field arithmetic. In contrast, SNARK-friendly hashes such as Poseidon, Rescue, or MiMC are designed to align with finite-field operations, reducing constraint counts and improving prover time. Some zkVMs also isolate hashing into specialized components or choose trace layouts that make hash-heavy workloads cheaper to prove.
3.1.3 Projects
Some of the prominent zkVMs over the last decade are TinyRAM [13]; Hawk [50]; Zinc [57]; RISC Zero [16], CairoVM [92], Leo [4], Miden [80], Triton [69]; OlaVM [91]; Powdr [81], Jolt [8]; SP1 [97], Nexus [56], and Valida [54]. Table 3 summarizes representative design choices that recur in the zkVM ecosystem like arithmetization target, proof family, hash strategy, and recursion support. Across many current systems, two recurring patterns appear: VM-style designs that prove an execution trace often favor transparent setups and hash-based commitments, frequently at the cost of larger proofs and hash-heavy traces; SNARK-style backends often emphasize smaller proofs and faster verification, but may rely on stronger setup assumptions depending on the concrete scheme. These tradeoffs matter directly for rollups, zkEVMs, and general verifiable compute.
Table 3: Representative zkVM design choices.
zkVM Execution encoding Proof family Hash strategy Recursion / composition
RISC Zero RISC-V trace + constraints STARK (optional SNARK wrap) hash-heavy (SHA-2 common) supported
CairoVM AIR (trace constraints) STARK hash-heavy trace supported
Miden VM AIR (stack trace) STARK ZK-friendly perm hash supported
A project in the space is RISC Zero’s zkVM [16] which allows users to demonstrate the accurate execution of arbitrary Rust and C++ code on an embedded RISC-V micro-processor. Starkware [92], the company behind zkSTARKs, developed CairoVM. CairoVM implements a STARK-based von-Neumann architecture, supporting traditional language features like conditional branching, function calls, recursion, and a restrictive memory model. Polygon’s Miden VM [80] is a zkVM designed for zero-knowledge rollups, which enhances Ethereum with features like parallel transaction execution and client-side proving.
3.2 Domain Specific Languages
Table 4: Comparison of zkDSLs